How do I figure out this problem? I have multiples like this, but need help understanding. Thank you.
if 480 million dollars was the total sales for products purchased online, how much of it was spent on apparel?
Ski products Purchased Online:
Lift tickets= 11/50
Accessories= 7/50
Apparel= 9/20
Equipment= 19/100
PLEASE HELP!!

Answer by richwmiller(17219)

9/20*480=
9*12=108 million